### v1

# Unconventional tunnel magnetoresistance scaling with altermagnets

Abstract
In conventional magnetic tunnel junctions (MTJs), the tunnel magnetoresistance (TMR) typically increases with barrier thickness as electron transmission in the antiparallel configuration decays faster than that of the parallel configuration. In this work, we reveal an anomalous scaling effect in altermagnetic tunnel junctions (AMTJs), where the TMR decreases anomalously with an increasing barrier thickness. The anomalous scaling originates from the overlapping spin-split branches form a transmission path that cannot be suppressed in the antiparallel state. Such phenomena is explained by a double-barrier model and is further demonstrated using ab initio quantum transport simulations in 2D V2Te2O/Cr2Se2O/V2Te2O-based AMTJ, where the TMR anomalously decreases from 220% to 40% as the layer number of Cr2Se2O increases from 1 to 5. Our work identifies a peculiar unexpected transport characteristic of AMTJ, providing a fundamental limit on AMTJ device design and illustrating the potential optimal design of AMTJ at the ultrascaled monolayer limit.
- Keywords
quantum transport simulation, altermagnetic, magnetic tunnel junction, scaling be-
havior, first-principles calculation
Introduction. Magnetic tunnel junctions (MTJs) are a fundamental component of spintronics, which constitute a nonvolatile memory function through the giant magnetoresistance effect [1, 2, 3, 4, 5, 6, 7, 8, 9]. The scaling of MTJs is a key trend in advancement of spintronics to achieve high-density storage, high-speed operation, and efficient spin orbital torque (SOT) [10, 11, 12, 13, 14, 15, 16]. Ferromagnetic memory faces difficulties to achieve higher storage density and device miniaturization due to the presence of stray fields [17, 18, 19, 20]. Antiferromagnetic tunnel junctions have gained attention for their robustness against magnetic perturbations, absence of stray fields, high Néel temperature, and ultrafast spin dynamics [21, 22, 23, 24, 25, 26]. However, conventional antiferromagnets exhibit degenerate energy bands, with only hidden spin polarization available for control. Altermagnets (AMs), featuring alternating spin polarizations in both real space and momentum space, have recently been recognized as a type of unconventional magnets [27, 28, 29, 30, 31, 32]. They exhibit a giant spin-splitting without a net magnetic moment, thus offering an ideal platform for ultracompact integration of MTJs [33].
Tunnel magnetoresistance (TMR), a key metric characterizing the performance of MTJs, is defined as
where and denote the conductance of the device in the parallel (P) and anti-parallel (AP) magnetic configurations, respectively [34]. In ferromagnetic and A-type antiferromagnetic tunnel junctions, TMR typically increases as the barrier layer thickness increases [35, 36, 37, 38, 39, 40, 41]. For example, in graphite/2D CrPS4/graphite MTJs, the TMR ratio rises by approximately when the number of CrPS4 layers is increased from 2 to 7 [42]. A similar trend is observed in Ag/CrI3/Ag MTJs, where the TMR ratio grows from 103% to 109% with the barrier thickness of CrI3 varying from 2 to 12 layers [37]. This behavior can be explained by the different decay rates of conductance in the P and AP configurations as described by the Julliere model [38, 43, 44, 34]. As shown in Fig. 1(a), the conductance is dominated by Bloch states with a small decay rate in P state, which is also called the majority spin conductance. In the AP configuration, the conductance is strongly suppressed due to spin mismatch between the electrodes, and it is mainly contributed by interface resonance states with large decay rate. As a result, TMR decreases as the barrier becomes thinner according to , where represents the difference in the decay rate between the AP and the P configurations.
Anomalous scaling in AMTJs. In this work, we reveal an anomalous scaling behavior in altermagnetic tunnel junctions (AMTJs) [Fig. 1(b)], whose TMR decreases as the barrier thickness increases. This anomalous scaling originates from the distinct morphology of the altermagnetic Fermi surface, where the momentum-space overlap between the spin-split branches forms a persistent transmission channel that cannot be suppressed in the AP configuration. The states that suppress the increase in TMR are, in fact, the spin-degenerate Bloch states in altermagnets. Such spin degeneracy is general in altermagnets and protected by the symmetry when the little group of momentum includes at least one real-space symmetry operation that connects the opposite-spin sublattices. [29, 33]. In contrast, ferromagnetic or A-type antiferromagnetic MTJs as discussed above possess fully spin-split bands without overlap, allowing efficient suppression of transmission in the AP configuration. To illustrate the anomalous scaling, we develop a double-barrier model, which sub-divides the spin-split Fermi surfaces into an outer region and a central region, labelled as A and B in Figs. 1(c) and 1(d), respectively. In the AP configuration, Neel vectors of the incident and transmission leads are misaligned. Electron transmission mediated by states in Region A can be effectively switched off due to spin-mismatch in this case. However, transport channel in Region B remains open even in the AP configuration. This arises because of the overlapping of two spin-split Fermi surfaces which create a spin-degenerate transmission highway that remains open in both P and AP states [Figs. 1(c) and 1(d)]. As the barrier thickness increases, electron tunneling through the spin-degenerate highway (provided by the electronic states in Region B) remains high with higher OFF state current. Consequently, TMR reduces with increasing barrier thickness, leading to the anomalous scaling behavior as compared to the opposite behavior in conventional MTJs.
To validate this mechanism, we investigate the scaling behavior of AMTJ composed of 2D V2Te2O/Cr2Se2O/V2Te2O using ab initio quantum transport simulation. By increasing the Cr2Se2O layer number from 1 to 5 layers, the TMR ratio decreases from 220% to 40%, which is in stark contrasts with the conventional scaling behavior observed in a wider of different MTJs based on different materials [35, 36, 37, 38, 39, 40, 41]. This work uncovers a previously unrecognized transport mechanism in AMs, emphasizing that spin-degenerate transmission channels can fundamentally limit the spin-selective transport efficiency in AMTJ. These findings reveal a potential limitation in the scaling behavior of AMTJ, highlighting the importance of controlling spin-degenerate channels to optimize TMR in AMTJ-based devices.
Anomalous scaling mechanism. According to the tunneling theory, electron transmission decays exponentially as where is the barrier thickness [45]. Here is the imaginary part of the wavevector (), which can be written as
where is the wavevector parallel to the tunneling interface, is the Fermi level [38, 43], and potential barrier is a function of and spin . We denote two different tunneling barriers for spin-degenerate transport mediated by states in Region B and spin-nondegenerate transport mediated by states in Region A as and , respectively. In AMTJ, the transport channel associated with is always open regardless of the P and AP configurations. In contrast, the transport channel through can be more effectively switched ON or OFF, thus playing the key role in determining the scaling of TMR. The total conductance under is composed of
where is quantum conductance ( and are elementary charge and Planck constant, respectively) and , and the index corresponds to the contribution of electron transmission mediated through Region A and B, respectively. The total conductance in the P configuration is . In the AP configuration, the tunneling through interface resonance states in Region B can be neglected, resulting in a conductance of . The -related TMR ratio is
The total TMR is the summation . Here we consider two scenarios: (i) If , the TMR increases with increasing barrier thickness , which is consistent with the conventional scaling behavior. (ii) If , which generally occurs in AMTJ, the conductance decays more rapidly than as increases as shown in Fig. 1(b). As a result, the TMR increases with decreasing , as shown at the bottom of Fig. 1(b), which corresponds to the anomalous scaling effect. We next demonstrate this anomalous scaling behavior in a representative AMTJ, corresponding to the second case discussed above.
First principle calculation. We calculate the scaling behavior of 2D AMTJ based on 2D AM metal V2Te2O (VTO) and AM semiconductor Cr2Se2O (CSO). VTO and CSO are both room-temperature AM with high Néel temperature of 740 K and 350 K, respectively. The lattice mismatch of VTO/CSO heterojunction is less than 0.6% [46, 47]. The atomic structure of these two materials exhibits , , , and symmetry [Fig. 2(a)]. The spin polarization of Cr or V atoms are opposite in the unit cell [bottom right in Fig. 2(a)]. With such magnetic ordering, symmetry (where represents a half-lattice translation along [110]) is broken, leading to unconventional spin splitting in reciprocal space. This phenomenon is further validated by the density functional theory (DFT)-calculated monolayer (ML) band structures [Figs. 2(b) and (d)]. The -resolved spin-splitting energy at the conduction band minimum (CBM) of CSO and VTO [Figs. 2(c) and (e)] indicate that they both possess -wave anisotropy with spin degeneration around point. This band structure features a spin-degenerate region near the point and spin-split regions away from it, corresponding to Region B and Region A in the aforementioned mechanism. It is worth noting that this condition is not limited to the present system; similar characteristics can also be found in altermagnets with -wave or -wave anisotropy. The CSO is serving as an insulating barrier with a bandgap of 0.97 eV. The CSO in the insulating layer is AB-stacked, and its structure is shown in Fig. S2 in Supplementary Material [48]. To ensure consistent spin splitting in -space across different layers, an interlayer out-of-plane antiferromagnetic (AFM) spin alignment is adopted within the same material, forming a G-type configuration [49]. We construct a tight-binding (TB) model to describe the spin splitting characteristics of the CSO, which is consistent with the results from DFT calculations (Fig. S1 in Supplementary Material [48]).
The decay of the tunneling electrons of bulk CSO is obtained by calculating complex band structure [50, 51]. Figure 3 shows the out-of-plane lowest decay rate of spin-up and spin-down electrons of CSO near the Fermi level. The decay characteristics in altermagnets exhibit an alternating nature. Due to spin degeneracy, spin-up and spin-down electrons near the point, i.e. in Region A, have nearly identical decay rates. In contrast, Region B corresponds to the spin-split states along the –X and –Y directions, where the spin polarization reverses between the two paths with the relation of . The transmitted wave decays more slowly in Region B than in Region A, indicating that the decay channel associated with Region B is more robust than that of Region A.
We consider a VTO/nL-CSO/VTO AMTJ composed of an n-layer CSO barrier sandwiched between layered-bulk VTO electrodes (Fig. 4a). The Néel vectors of both the electrodes and the barrier can reorient independently, giving rise to three distinct magnetic configurations: the parallel (P) configuration, with all Néel vectors aligned (Fig. 4b); the intermediate (I) configuration, where the barrier is antiparallel to the identically aligned electrodes (Fig. 4c); and the antiparallel (AP) configuration, where the two electrodes have opposite Néel vectors (Fig. 4d) (see Fig. S3 for detailed spin alignments for these configurations [48]).
We perform DFT calculations coupled with the nonequilibrium Green’s function (NEGF) method, as implemented in QuantumATK (see Supplementary Materials for computational details [48]) [52], to assess the transport properties of the AMTJ. In the VTO/1L-CSO/VTO AMTJ, switching from the AP state yields TMR ratios of about 220% for the P state and 137% for the I state. The distinct transport behaviors between P and I configurations can be attributed to the spin-dependent tunneling barrier. The difference in transmission between the P and I configurations arises from the lower tunneling barrier in the P state, where the spin-splitting of the electrodes and barrier are mutually aligned (Figs. 4e,f). Likewise, the transmission eigenstates for different spin channels further reveal that the tunneling barrier due to the spin-splitting mismatch between CSO and VTO effectively suppresses electron transmission in the I state (see Figs. S4 in Supplementary Material [48]). The lowest transmission observed in the AP state (Fig. 4g) indicates that spin matching between electrodes plays a more dominant role in tunneling than spin alignment within the barrier. Especially, spin-degenerate states (near the point) consistently exhibit lower tunneling barriers due to the constant spin alignment in all three magnetic configurations and dominate the transport in the AP state.
Next, we investigate the tunneling of AMTJ with varying CSO layer number. With increasing CSO thickness, all three magnetic configurations exhibit exponential decay in transmission. However, the decay rates differ significantly: the AP configuration shows the slowest decay (only 31%), in contrast to 72.7% and 59.3% reductions in the P and I states, respectively. This trend highlights a persistent transmission channel in the AP configuration that is less sensitive to barrier thickness. As a result, the TMR drops anomalously—from 220% to 40% and from 137% to 39% for P-to-AP and I-to-AP switching, respectively—as the number of CSO layers increases from 1 to 5 (Fig. 5a). The detailed energy-dependent transmissions for different magnetic configurations are shown in Fig. S3(b)–(c) (Supplementary Material [48]).
In summary, we reveal an anomalous scaling behavior of decreasing TMR with increasing barrier thickness in AMTJ, which deviates from conventional expectations. The underlying mechanism of this anomalous scaling lies in the presence of a persistent transmission channel arising from spin-degenerate states, which remains largely unaffected by switching between the P and AP configurations. To qualitatively interpret this behavior, we further construct a simplified double-barrier model that distinguishes between effective and ineffective spin-selective transmission. The effective part, originating from spin-split states, decays rapidly with increasing barrier thickness. While the ineffective part exhibits weak decay and ultimately dominates the overall transmission at larger thicknesses. This contrasts with conventional ferromagnetic MTJs, where the bandstructure is fully split and only the effective spin-selective transmission is present. To validate this mechanism, we conduct ab initio quantum transport simulations on a V2Te2O/Cr2Se2O/V2Te2O AMTJ, demonstrating that increasing the Cr2Se2O layer thickness from 1 to 5 layers led to a significant TMR degradation from 220% to 40%. Our results not only provide new insights into the fundamental transport behavior of altermagnets but also highlight a key design consideration fundamentally impacting the performance of magnetic tunneling junction based on altermagnets.
Experimental implementation. We note that bulk materials with similar structure to these two materials, KV2Se2O and SbV2Te2O, have been successfully synthesized in recent experiments [53, 54]. Spin- and angle-resolved photoemission spectroscopy (SARPES) measurements confirm them as -wave anisotropic AM. Moreover, its 2D form can be obtained via topochemical deintercalation [55], thus suggesting V2Te2O and Cr2Se2O as potential feasible material platform for experimental realization.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China (No. 12274002 and 91964101), the Ministry of Science and Technology of China (No. 2022YFA1203904), the Fundamental Research Funds for the Central Universities, the High-performance Computing Platform of Peking University and the MatCloud + high-throughput materials simulation engine, and the National Research Foundation, Singapore, under its Frontier Competitive Research Programme (NRF-F-CRP-2024-0001). L. K. A. is supported by the Singapore ASTAR IRG (M23M6c0102).References
- Han et al. [2024] L. Han, X. Fu, R. Peng, X. Cheng, J. Dai, L. Liu, Y. Li, Y. Zhang, W. Zhu, H. Bai, Y. Zhou, S. Liang, C. Chen, Q. Wang, X. Chen, L. Yang, Y. Zhang, C. Song, J. Liu, and F. Pan, Electrical 180° switching of Néel vector in spin-splitting antiferromagnet, Sci. Adv. 10, eadn0479 (2024).
- Dieny et al. [2020] B. Dieny, I. L. Prejbeanu, K. Garello, P. Gambardella, P. Freitas, R. Lehndorff, W. Raberg, U. Ebels, S. O. Demokritov, J. Akerman, A. Deac, P. Pirro, C. Adelmann, A. Anane, A. V. Chumak, A. Hirohata, S. Mangin, S. O. Valenzuela, M. C. Onbaşlı, M. d’Aquino, G. Prenat, G. Finocchio, L. Lopez-Diaz, R. Chantrell, O. Chubykalo-Fesenko, and P. Bortolotti, Opportunities and challenges for spintronics in the microelectronics industry, Nat. Electron. 3, 446 (2020).
- Schwierz et al. [2015] F. Schwierz, J. Pezoldt, and R. Granzner, Two-dimensional materials and their prospects in transistor electronics, Nanoscale 7, 8261 (2015).
- Du et al. [2023] A. Du, D. Zhu, K. Cao, Z. Zhang, Z. Guo, K. Shi, D. Xiong, R. Xiao, W. Cai, J. Yin, S. Lu, C. Zhang, Y. Zhang, S. Luo, A. Fert, and W. Zhao, Electrical manipulation and detection of antiferromagnetism in magnetic tunnel junctions, Nat. Electron. 6, 425 (2023).
- Hirohata et al. [2020] A. Hirohata, K. Yamada, Y. Nakatani, I.-L. Prejbeanu, B. Diény, P. Pirro, and B. Hillebrands, Review on spintronics: Principles and device applications, J. Magn. Magn. Mater. 509, 166711 (2020).
- Zhang et al. [2024] B. Zhang, P. Lu, R. Tabrizian, P. X.-L. Feng, and Y. Wu, 2D magnetic heterostructures: spintronics and quantum future, Npj Spintron. 2, 6 (2024).
- Bhatti et al. [2017] S. Bhatti, R. Sbiaa, A. Hirohata, H. Ohno, S. Fukami, and S. N. Piramanayagam, Spintronics based random access memory: a review, Mater. Today 20, 530 (2017).
- Zhang et al. [2021] L. Zhang, J. Zhou, H. Li, L. Shen, and Y. P. Feng, Recent progress and challenges in magnetic tunnel junctions with 2D materials for spintronic applications, Appl. Phys. Rev. 8, 021308 (2021).
- Wu and Ang [2022] Q. Wu and L. K. Ang, Giant tunneling magnetoresistance in atomically thin VSi2N4/MoSi2N4/VSi2N4 magnetic tunnel junction, Appl. Phys. Lett. 120, 022401 (2022).
- Baltz et al. [2018] V. Baltz, A. Manchon, M. Tsoi, T. Moriyama, T. Ono, and Y. Tserkovnyak, Antiferromagnetic spintronics, Rev. Mod. Phys. 90, 015005 (2018).
- Jungwirth et al. [2016] T. Jungwirth, X. Marti, P. Wadley, and J. Wunderlich, Antiferromagnetic spintronics, Nat. Nanotechnol. 11, 231 (2016).
- Han et al. [2021] B. Han, B. Zhang, S. Sun, B. Wang, Y. Guo, and J. Cao, The thickness dependence of the field-like spin–orbit torque in heavy metal/CoFeB/MgO heterostructures, J. Appl. Phys. 130, 213902 (2021).
- Chen et al. [2024] H. Chen, G. Zhou, H. Ji, Q. Qin, S. Shi, Q. Shen, P. Yao, Y. Cao, J. Chen, Y. Liu, H. Wang, W. Lin, Y. Yang, J. Jia, X. Xu, J. Chen, and L. Liu, Spin-Orbit Torque Switching of Magnetization in Ultra-Thick Ferromagnetic Layers, Adv. Funct. Mater. 34, 2403107 (2024).
- Choi et al. [2022] G. Choi, J. Ryu, R. Thompson, J.-G. Choi, J. Jeong, S. Lee, M.-G. Kang, M. Kohda, J. Nitta, and B.-G. Park, Thickness dependence of spin–orbit torques in Pt/Co structures on epitaxial substrates, APL Mater. 10, 011105 (2022).
- Lee et al. [2021] M. H. Lee, G. Go, Y. J. Kim, I. H. Cha, G. W. Kim, T. Kim, K.-J. Lee, and Y. K. Kim, Spin–orbit torques in normal metal/Nb/ferromagnet heterostructures, Sci. Rep. 11, 21081 (2021).
- Manchon et al. [2019] A. Manchon, J. Železný, I. Miron, T. Jungwirth, J. Sinova, A. Thiaville, K. Garello, and P. Gambardella, Current-induced spin-orbit torques in ferromagnetic and antiferromagnetic systems, Rev. Mod. Phys. 91, 035004 (2019).
- sys [2024] Spintronics for ultra-low-power circuits and systems, Nat. Rev. Electr. Eng. 1, 691 (2024).
- Masuda et al. [2024] H. Masuda, T. Seki, J.-i. Ohe, Y. Nii, H. Masuda, K. Takanashi, and Y. Onose, Room temperature chirality switching and detection in a helimagnetic MnAu2 thin film, Nat. Commun. 15, 1999 (2024).
- Han et al. [2023] J. Han, R. Cheng, L. Liu, H. Ohno, and S. Fukami, Coherent antiferromagnetic spintronics, Nat. Mater. 22, 684 (2023).
- Xiong et al. [2022] D. Xiong, Y. Jiang, K. Shi, A. Du, Y. Yao, Z. Guo, D. Zhu, K. Cao, S. Peng, W. Cai, D. Zhu, and W. Zhao, Antiferromagnetic spintronics: An overview and outlook, Fundam. Res. 2, 522 (2022).
- Qin et al. [2023] P. Qin, H. Yan, X. Wang, H. Chen, Z. Meng, J. Dong, M. Zhu, J. Cai, Z. Feng, X. Zhou, L. Liu, T. Zhang, Z. Zeng, J. Zhang, C. Jiang, and Z. Liu, Room-temperature magnetoresistance in an all-antiferromagnetic tunnel junction, Nature 613, 485 (2023).
- Shao and Tsymbal [2024] D.-F. Shao and E. Y. Tsymbal, Antiferromagnetic tunnel junctions for spintronics, Npj Spintron. 2, 1 (2024).
- Olsen [2024] T. Olsen, Antiferromagnetism in two-dimensional materials: progress and computational challenges, 2D Mater. 11, 033005 (2024).
- Shao et al. [2023] D.-F. Shao, Y.-Y. Jiang, J. Ding, S.-H. Zhang, Z.-A. Wang, R.-C. Xiao, G. Gurung, W. Lu, Y. Sun, and E. Y. Tsymbal, Néel Spin Currents in Antiferromagnets, Phys. Rev. Lett. 130, 216702 (2023).
- Gu et al. [2023] P. Gu, C. Wang, D. Su, Z. Dong, Q. Wang, Z. Han, K. Watanabe, T. Taniguchi, W. Ji, Y. Sun, and Y. Ye, Multi-state data storage in a two-dimensional stripy antiferromagnet implemented by magnetoelectric effect, Nat. Commun. 14, 3221 (2023).
- Chen et al. [2023] X. Chen, T. Higo, K. Tanaka, T. Nomoto, H. Tsai, H. Idzuchi, M. Shiga, S. Sakamoto, R. Ando, H. Kosaki, T. Matsuo, D. Nishio-Hamane, R. Arita, S. Miwa, and S. Nakatsuji, Octupole-driven magnetoresistance in an antiferromagnetic tunnel junction, Nature 613, 490 (2023).
- Šmejkal et al. [2022a] L. Šmejkal, J. Sinova, and T. Jungwirth, Beyond Conventional Ferromagnetism and Antiferromagnetism: A Phase with Nonrelativistic Spin and Crystal Rotation Symmetry, Phys. Rev. X 12, 031042 (2022a).
- Fender et al. [2025] S. S. Fender, O. Gonzalez, and D. K. Bediako, Altermagnetism: A Chemical Perspective, J. Am. Chem. Soc. 147, 2257 (2025).
- Šmejkal et al. [2022b] L. Šmejkal, J. Sinova, and T. Jungwirth, Emerging Research Landscape of Altermagnetism, Phys. Rev. X 12, 040501 (2022b).
- Liu et al. [2025] Q. Liu, X. Dai, and S. Blügel, Different facets of unconventional magnetism, Nat. Phys. 21, 329 (2025).
- Krempaský et al. [2024] J. Krempaský, L. Šmejkal, S. W. D’Souza, M. Hajlaoui, G. Springholz, K. Uhlířová, F. Alarab, P. C. Constantinou, V. Strocov, D. Usanov, W. R. Pudelko, R. González-Hernández, A. Birk Hellenes, Z. Jansa, H. Reichlová, Z. Šobáň, R. D. Gonzalez Betancourt, P. Wadley, J. Sinova, D. Kriegner, J. Minár, J. H. Dil, and T. Jungwirth, Altermagnetic lifting of Kramers spin degeneracy, Nature 626, 517 (2024).
- Song et al. [2025] C. Song, H. Bai, Z. Zhou, L. Han, H. Reichlova, J. H. Dil, J. Liu, X. Chen, and F. Pan, Altermagnets as a new class of functional materials, Nat. Rev. Mater. 1 (2025).
- Bai et al. [2024] L. Bai, W. Feng, S. Liu, L. Šmejkal, Y. Mokrousov, and Y. Yao, Altermagnetism: Exploring New Frontiers in Magnetism and Spintronics, Adv. Funct. Mater. , 2409327 (2024).
- Tsymbal et al. [2003] E. Y. Tsymbal, O. N. Mryasov, and P. R. LeClair, Spin-dependent tunnelling in magnetic tunnel junctions, J. Phys. Condens. Matter 15, R109 (2003).
- Yuasa et al. [2004] S. Yuasa, T. Nagahama, A. Fukushima, Y. Suzuki, and K. Ando, Giant room-temperature magnetoresistance in single-crystal Fe/MgO/Fe magnetic tunnel junctions, Nat. Mater. 3, 868 (2004).
- Yang et al. [2024] J. Yang, B. Wu, S. Zhao, S. Liu, J. Lu, S. Li, and J. Yang, Multistate magnetic tunnel junction based on a single two-dimensional van der waals antiferromagnet, Phys. Rev. Appl. 22, 014017 (2024).
- Wu et al. [2022] B. Wu, J. Yang, R. Quhe, S. Liu, C. Yang, Q. Li, J. Ma, Y. Peng, S. Fang, J. Shi, J. Yang, J. Lu, and H. Du, Scaling Behavior of Magnetoresistance with the Layer Number in CrI3 Magnetic Tunnel Junctions, Phys. Rev. Appl. 17, 034030 (2022).
- Butler et al. [2001] W. H. Butler, X.-G. Zhang, T. C. Schulthess, and J. M. MacLaren, Spin-dependent tunneling conductance of Fe | MgO | Fe sandwiches, Phys. Rev. B 63, 054416 (2001).
- Chi et al. [2025] B. Chi, L. Jiang, Y. Zhu, G. Yu, C. Wan, and X. Han, Anisotropic spin filtering by an altermagnetic barrier in magnetic tunnel junctions, Phys. Rev. Appl. 23, 014013 (2025).
- Emoto et al. [2024] S. Emoto, H. Kusunose, Y.-C. Lin, H. Sun, S. Masuda, S. Fukamachi, K. Suenaga, T. Kimura, and H. Ago, Synthesis of Few-Layer Hexagonal Boron Nitride for Magnetic Tunnel Junction Application, ACS Appl. Mater. Interfaces 16, 31457 (2024).
- Zhou et al. [2020] M. Zhou, H. Jin, and Y. Xing, In-Plane Dual-Gated Spin-Valve Device Based on the Zigzag Graphene Nanoribbon, Phys. Rev. Appl. 13, 044006 (2020).
- Yang et al. [2021] J. Yang, S. Fang, Y. Peng, S. Liu, B. Wu, R. Quhe, S. Ding, C. Yang, J. Ma, B. Shi, L. Xu, X. Sun, G. Tian, C. Wang, J. Shi, J. Lu, and J. Yang, Layer-Dependent Giant Magnetoresistance in Two-Dimensional CrPS4 Magnetic Tunnel Junctions, Phys. Rev. Appl. 16, 024011 (2021).
- MacLaren et al. [1997] J. M. MacLaren, X.-G. Zhang, and W. H. Butler, Validity of the Julliere model of spin-dependent tunneling, Phys. Rev. B 56, 11827 (1997).
- Julliere [1975] M. Julliere, Tunneling between ferromagnetic films, Phys. Lett. A 54, 225 (1975).
- Band and Avishai [2013] Y. B. Band and Y. Avishai, Quantum mechanics with applications to nanotechnology and information science, Academic Press (2013).
- Qirui et al. [2023] C. Qirui, Z. Yingmei, Y. Xiong, C. Ping, and Y. Hongxin, Giant spin-hall and tunneling magnetoresistance effects based on a two-dimensional nonrelativistic antiferromagnetic metal, Phys. Rev. B 108, 024410 (2023).
- Gong et al. [2024] J. Gong, Y. Wang, Y. Han, Z. Cheng, X. Wang, Z. Yu, and Y. Yao, Hidden Real Topology and Unusual Magnetoelectric Responses in Two‐Dimensional Antiferromagnets, Adv. Mater. 36, 2402232 (2024).
- [48] S. Material, Supplementary material, .
- Peng et al. [2025] R. Peng, J. Yang, L. Hu, W.-L. Ong, P. Ho, C. Siong Lau, J. Liu, and Y. Sin Ang, All-electrical layer-spintronics in altermagnetic bilayers, Mater. Horiz. (2025).
- Mavropoulos et al. [2000] P. Mavropoulos, N. Papanikolaou, and P. H. Dederichs, Complex Band Structure and Tunneling through Ferromagnet/Insulator/Ferromagnet Junctions, Phys. Rev. Lett. 85, 1088 (2000).
- Lukashev et al. [2012] P. V. Lukashev, A. L. Wysocki, J. P. Velev, M. van Schilfgaarde, S. S. Jaswal, K. D. Belashchenko, and E. Y. Tsymbal, Spin filtering with EuO: Insight from the complex band structure, Phys. Rev. B 85, 224414 (2012).
- Smidstrup et al. [2020] S. Smidstrup, T. Markussen, P. Vancraeyveld, J. Wellendorff, J. Schneider, T. Gunst, B. Verstichel, D. Stradi, P. A. Khomyakov, U. G. Vej-Hansen, M.-E. Lee, S. T. Chill, F. Rasmussen, G. Penazzi, F. Corsetti, A. Ojanperä, K. Jensen, M. L. N. Palsgaard, U. Martinez, A. Blom, M. Brandbyge, and K. Stokbro, Quantumatk: an integrated platform of electronic and atomic-scale modelling tools, J. Phys. Condens. Matter 32, 015901 (2020).
- Jiang et al. [2025] B. Jiang, M. Hu, J. Bai, Z. Song, C. Mu, G. Qu, W. Li, W. Zhu, H. Pi, Z. Wei, Y.-J. Sun, Y. Huang, X. Zheng, Y. Peng, L. He, S. Li, J. Luo, Z. Li, G. Chen, H. Li, H. Weng, and T. Qian, A metallic room-temperature d-wave altermagnet, Nat. Phys. (2025).
- Zhang et al. [2025] F. Zhang, X. Cheng, Z. Yin, C. Liu, L. Deng, Y. Qiao, Z. Shi, S. Zhang, J. Lin, Z. Liu, M. Ye, Y. Huang, X. Meng, C. Zhang, T. Okuda, K. Shimada, S. Cui, Y. Zhao, G.-H. Cao, S. Qiao, J. Liu, and C. Chen, Crystal-symmetry-paired spin–valley locking in a layered room-temperature metallic altermagnet candidate, Nat. Phys. (2025).
- Lin et al. [2018] H. Lin, J. Si, X. Zhu, K. Cai, H. Li, L. Kong, X. Yu, and H.-H. Wen, Structure and physical properties of CsV2Se2-xO and V2Se2O, Phys. Rev. B 98, 075132 (2018).